# Probabilistic Analysis for Scalogram Ridges in Signal Processing

### v2

@@ 6 Conclusion
In this work, we focus on providing a systematic study of Steps I and II as outlined in the Introduction. In addition to proposing a mathematical framework and definition for ridge analysis, we establish key properties of the ridge when viewed as a random process. Although we do not address Steps III and IV in detail, we briefly mention that most existing methods aim to fit a curve that captures the maximal energy on the TFR of $Y$, using various weighting and fitting strategies [carmona1997characterization, carmona1999multiridge, Ozkurt_Savaci_2005, iatsenko2016extraction, meignen2017demodulation, Zhu_Zhang_Gao_Li_2019, Legros_Fourer_2021]. When multiple oscillatory components are present, the IF separation condition has been employed to support algorithm design [Laurent_Meignen_2021]. In cases involving non-sinusoidal oscillations, geometric structures in the TFR can be leveraged to improve extraction accuracy [su2024ridge], and chirp characteristics can also be incorporated to enhance performance [Colominas_Meignen_Pham_2020]. Moreover, specialized algorithms have been proposed to handle scenarios with significant spectral interference [meignen2023new].
@@
Moreover, in real-world applications, the oscillatory pattern of a signal may change abruptly, rendering the considered AHM unsuitable. In such cases, the associated ridge becomes discontinuous, and ridge detection must account for change points. Additionally, noise in practice is often nonstationary, further complicating the problem. Extending the current framework to handle these challenges is both difficult and important. We leave further investigation of these directions to future work.
@@ 7.1 Proof of Theorem 1
Theorem 1(b) follows from the fact that the distribution of half the sum of two independent chi-square random variables, each with one degree of freedom, is an exponential distribution with a mean of one.
@@
where $\int^{{}^{\prime}}$ means that the integral excludes the diagonal hyperplanes $\lambda_{1}=\mp\lambda_{2}$ [major1981lecture]. Similarly,
@@
By the product formula for the second Wiener chaos [major1981lecture, nourdin2012normal],
@@
The proof of the first equality in Theorem 1(c) is concluded by substituting (52) into (49).
@@
The second equality in Theorem 1(c) implies that
@@
The proof of the second equality in Theorem 1(c) is concluded by noting that
@@ 7.2 Proof of Proposition 1
Step 1. To prove that the sample paths of the Gaussian field $W_{\Phi}$ are almost surely continuous on $\mathbb{R}\times(0,\infty)$, according to [dudley1973sample, p.73 and p.93], it suffices to show the existence of a constant $q>0$ such that for any $\delta>0$, there exists a constant $C_{\delta}$ satisfying
@@
To prove that sample paths of $W_{\Phi}$ are continuously differentiable almost surely, according to [potthoff2010sample], it suffices to show that (a) the difference quotient of $W_{\Phi}$ along the time direction converges to $\partial_{t}W_{\Phi}$ in quadratic mean, (b) the difference quotient of $W_{\Phi}$ along the scale direction converges to $\partial_{s}W_{\Phi}$ in quadratic mean, and (c) both $\partial_{t}W_{\Phi}$ and $\partial_{s}W_{\Phi}$ are continuous in quadratic mean and almost surely have continuous realizations. We verify these conditions as follows.
@@
Proof of (c): To prove that $\partial_{t}W_{\Phi}$ and $\partial_{s}W_{\Phi}$ not only are continuous in the quadratic mean but also almost surely have modification that is sample continuous, according to [dudley1973sample, p.73 and p.93] (see also [potthoff2009sample, Corollary 4.6]), it suffices to show that for any $\delta>0$, there exists a constant $C_{\delta}$ such that
@@ 7.3 Proof of Corollary 1
Proof of (11): Based on the continuity of the sample paths of $W_{\Phi}$ on $\mathbb{R}\times(0,\infty)$, it suffices to show that the sample paths of the Gaussian field $W_{\Phi}$ are almost surely continuous at the boundary $\mathbb{R}\times\{0\}$ if we define $W_{\Phi}(t,0)=0$ for any $t\in\mathbb{R}$. According to [dudley1973sample, p.73 and p.93], it suffices to show the existence of a constant $q>0$ such that there exists a constant $C$ satisfying
@@ 7.4 Proof of Theorem 2
where the second equality stems from Theorem 1(c). Therefore, for any $s>0$,
@@
where $\|\mathbf{z}\|^{2}=\mathbf{z}^{\top}\mathbf{z}$. By the result $\textup{Cov}\left(S_{\Phi}(s_{1}),S_{\Phi}(s_{2})\right)\geq 0$ given in Theorem 1(c), the Cauchy-Schwarz inequality, the condition in Assumption 1 that the spectral density $p$ is positive almost everywhere, and (70), we have
@@
    For $s$ satisfying $\textup{Cov}(S_{\Phi}(s),S_{\Phi}(s_{1}))=0$, by Theorem 1(c), we have $\mathbb{E}\left[W^{R}_{\Phi}(s)W^{R}_{\Phi}(s_{1})\right]=0$ and $\mathbb{E}\left[W^{R}_{\Phi}(s)W^{I}_{\Phi}(s_{1})\right]=0,$ which imply that $\Gamma(s,s_{1})=\mathbf{0}_{2\times 2}$ and $Q(\mathbf{z};s)=\mathbf{W}_{\perp}^{\top}(s)\mathbf{W}_{\perp}(s)$ for any $\mathbf{z}\in\mathbb{R}^{2}$. This indicates that $\{(\mathbf{z},Q(\mathbf{z},s))\mid\mathbf{z}\in\mathbb{R}^{2}\}$ is a horizontal plane.
By the fact that the supremum of a family of convex functions is still convex, both $M_{1}$ and $M_{2}$ are convex functions defined on $\mathbb{R}^{2}$. According to Alexandrov’s theorem [rockafellar1999second, Theorem 2.1], $M_{1}$ and $M_{2}$ have the second derivatives almost everywhere. More precisely, there exists a set $D\subset\mathbb{R}^{2}$, whose complement has Lebesgue measure zero, such that for $k\in\{1,2\}$ and every point $\mathbf{z}\in D$, there is a quadratic expansion in the form
@@
where $\mathbf{v}_{k}(\mathbf{z})$ is the gradient, and $\mathbf{H}_{k}(\mathbf{z})$ is the Hessian matrix of $M_{k}$ at $\mathbf{z}$. The function $o(\|\mathbf{w}\|^{2})$ satisfies $o(\|\mathbf{w}\|^{2})\|\mathbf{w}\|^{-2}\rightarrow 0$ as $\|\mathbf{w}\|\rightarrow 0$. By (73) (see also [rockafellar2009variational, Example 13.10]),
@@ 7.5 Proof of Theorem 3
The continuity of $S_{Y}(t,s;\omega)$ and the compactness of $0,\mathcal{S}(\omega)]$ allow us to apply Berge’s maximum theorem [hu1997handbook, Theorem 3.4] (see also [aliprantis2006infinite, p. 570] or Lemma 5 in the appendix), implying that the set-valued function $s_{Y}(\cdot;\omega)$ is upper hemicontinuous at every time point $t_{0}\in\mathcal{M}$. ∎
@@ 7.6 Proof of Corollary 2
Using the notation introduced at the beginning of Section 7.5, Theorems 2 and 3 show that there exists a subset $\widetilde{\Omega}$ of $\Omega$ with $\mathbb{P}(\widetilde{\Omega})=1$ such that for any $\omega\in\widetilde{\Omega}$,
@@ 7.7 Proof of Theorem 4
Because the wavelet $\psi$ is assumed to be analytic, $\mathbf{J}$ is a circularly symmetric complex Gaussian random vector, as shown in Theorem 1(a). By Lemma 1, for any $u>\mathbb{E}[g(\mathbf{J})]$,
@@ 7.9 Proof of Lemma 1
By the Gaussian concentration inequality [kuhn2023maximal, Corollary 8.5], for any $u>\mathbb{E}[g_{2d}(\mathbf{A}_{2d}\mathbf{N}_{2d})]$,
@@ 7.10 Proof of Lemma 2
Under Assumption 3(b1), $p(\lambda)=L(|\lambda|^{-1})|\lambda|^{H-1}$, where $L$ is slowly varying at infinity. According to [bingham1989regular, Theorem 1.5.3], for any slowly varying function $L$ and any $\delta>0$, $L(s)s^{-\delta}\rightarrow 0$ as $s\rightarrow\infty$. Thus, (113) implies that
@@ 7.11 Proof of Lemma 3
The proof idea can be traced back to the chaining argument in [talagrand2014upper]. We firstly create a sequence of sets, denoted as $\{\pi_{j}\}_{j\in\mathbb{N}\cup\{0\}}$, of finite points in $0,\infty)$ such that for each $j\in\mathbb{N}\cup\{0\}$,
@@
By Theorem 1(b), the random variable
@@ 7.12 Proof of Lemma 4
Interpretation of Theorem 5 by the single component case $M=1$: In the following, we omit the component-specific index $m$. That is, the symbols $A_{m}$, $L_{m}$, $\mu_{m,1}$$\mu_{m,4}$, and $\sigma_{m,1}$$\sigma_{m,4}$ are replaced by $A$, $L$, $\mu_{1}$$\mu_{4}$, and $\sigma_{1}$$\sigma_{4}$, respectively, to reflect the single-component scenario.
  • 1.
    $L$: A larger value of the second derivative of $|W_{f}(t,\cdot)|^{2}$ with respect to the scale variable at $s_{f}(t)$ indicates that $|W_{f}(t,\cdot)|$ exhibits a sharper peak at $s_{f}(t)$. This can be associated with a rapid change in the spectral content of the signal $f$. Hence, when $L$ is larger, the ridge point of $f$ is less likely to be disturbed. If the magnitude function $A$ varies slowly enough to allow the approximation
    $S_{f}(t,s)\approx A^{2}(t)S_{f/A}(t,s),$ (7.144)
    then $L$ is approximately proportional to $A^{2}(t)$. In this case, a larger $A$ similarly reduces the possibility of ridge point disturbance.
  • 2.
    $\mu_{1}$, $\mu_{2}$, $\mu_{3}$, $\mu_{4}$: Under the approximation (7.144),
    $\frac{6}{L}\max\left\{\mu_{1},\mu_{2},\mu_{3}^{2},\mu_{4}^{2}\right\}\approx\max\left\{\mathcal{O}\left(\frac{\sqrt{\textup{Var}(\Phi)}}{A}\right),\mathcal{O}\left(\frac{\textup{Var}(\Phi)}{A^{2}}\right)\right\}.$ (7.145)
  • 3.
    For any fixed $\varepsilon$ satisfying (28), (7.145) shows that the exponents
  • 4.
    and
    This also implies that, for any fixed $\varepsilon$ satisfying (28), the probability that $s_{Y}(t)$ deviates from $s_{f}(t)$ by more than $\varepsilon$ decreases exponentially fast when the ratio $A/\sqrt{\textup{Var}(\Phi)}$ increases.
@@
[MR1039321, Lemma (4.3.16)] If $\mathbb{X}$ is a topological space, $\mathbb{Y}$ a complete metric space and $r:\mathbb{D}\mapsto\mathbb{Y}$ a continuous mapping defined on a dense subset $\mathbb{D}$ of the space $\mathbb{X}$, then the mapping $r$ is extendable to a continuous mapping defined on the set consisting of all points of $\mathbb{X}$ at which the oscillation of $r$ is equal to zero. Here, we say that the oscillation of the mapping $r$ at a point $x\in\mathbb{X}$ is equal to zero if for every $\varepsilon>0$ there exists a neighborhood $U$ of the point $x$ such that

### v3

@@ 6 Conclusion
In this work, we focus on providing a systematic study of Steps I and II as outlined in the Introduction. In addition to proposing a mathematical framework and definition for ridge analysis, we establish key properties of the ridge when viewed as a random process. Although we do not address Steps III and IV in detail, we briefly mention that most existing methods aim to fit a curve that captures the maximal energy on the TFR, using various weighting and fitting strategies [6, 5, 35, 22, 33, 48, 25]. When multiple oscillatory components are present, the IF separation condition has been employed to support algorithm design [24]. In cases involving non-sinusoidal oscillations, geometric structures in the TFR can be leveraged to improve extraction accuracy [44], and chirp characteristics can also be incorporated to enhance performance [9]. Moreover, specialized algorithms have been proposed to handle scenarios with significant spectral interference [31].
@@
Finally, we comment on the scope of applicability and limitations of the proposed framework. The second half of this work focuses on the deviation of ridge points of a noisy signal from those of the corresponding noise-free signal. In other words, our analysis concerns the difference between the local maximizers of the scalograms of the noise-affected and noise-free signals, rather than the difference between the ridge points of the noisy signal and the IFs of the noise-free signal. As a result, the findings in this work can be extended to more general signal models, provided that the scalogram of the noise-free signal $f$ exhibits well-separated local maxima at each time. We shall mention that our current result is not sufficient to fully describe the general model like the adaptive non-harmonic model [47], defined as
$f_{w}(t)=\sum_{m=1}^{M}A_{m}(t)s_{m}\left(\phi_{m}(t)\right),$
where $M$, $A_{m}(t)$, $\phi_{m}(t)$, and $\phi_{m}^{\prime}(t)$ have the same interpretation as those of AHM, and $s_{m}$ is the wave-shape function of the $m$-th component, which is a $1$-periodic smooth function describing the non-sinusoidal oscillatory shape. Due to the non-sinusoidal oscillatory nature of $s_{m}$, $f_{w}$ can be written as
$f_{w}(t)=\sum_{m=1}^{M}\sum_{k=1}^{D_{m}}A_{m}(t)a_{m,k}\cos(2\pi k\phi_{m}(t)+\beta_{m,k}),$
where $D_{m}\geq 1$, $\beta_{m,k}\in[0,2\pi)$, and $\alpha_{m,k}\geq 0$ are the Fourier coefficients of $s_{m}$. With this expression, $f_{w}$ can be viewed as a generalized AHM with many IMT functions, while different IMT functions might have close or even crossover IFs. Since our results rely on the well-separated IFs of all IMT components, the analysis applies only to the special situation where $D_{m}$ is sufficiently small and the multiples of distinct $\phi_{m}^{\prime}(t)$ are well separated. Extending the theory to settings with spectral interference; that is, when IFs are close or intersect, remains an open problem for future ridge analysis.
Regarding noise models, our deviation bounds rely on Gaussian concentration inequalities for complex-valued Gaussian random variables arising from the analytic wavelet transform. These tools inherently require a Gaussian assumption. Extending the analysis to non-Gaussian noise would necessitate new maximal inequalities for complex-valued transforms of non-Gaussian random processes. For non-stationary noise, the covariance structure of the analytic wavelet transform becomes more involved, and corresponding theoretical extensions require further technical development. These issues are directly relevant in practice, where signals are discretized, and noise is typically non-Gaussian. Under mild non-stationarity assumptions (including controlled dependence and moments) and uniform sampling, it has been shown in [28] that the distribution of the discretized AWT of non-Gaussian, causal, non-stationary time series can be approximated by that of a suitable Gaussian sequence, so our results remain applicable. However, the case of non-uniform sampling remains open. Moreover, in real-world applications, oscillatory patterns may change abruptly, making the assumed adaptive harmonic model unsuitable. In such situations, ridges may become discontinuous, and ridge detection must incorporate change-point structure. Extending the current framework to address these challenges is both difficult and important, and we leave these directions for future work.
@@ 7.1 Proof of Proposition 1
Proposition 1(b) follows from the fact that the distribution of half the sum of two independent chi-square random variables, each with one degree of freedom, is an exponential distribution with a mean of one.
@@
where $\int^{{}^{\prime}}$ means that the integral excludes the diagonal hyperplanes $\lambda_{1}=\mp\lambda_{2}$ [29]. Similarly,
@@
By the product formula for the second Wiener chaos [29, 34],
@@
The proof of the first equality in Proposition 1(c) is concluded by substituting (54) into (51).
@@
The second equality in Proposition 1(c) implies that
@@
The proof of the second equality in Proposition 1(c) is concluded by noting that
@@ 7.2 Proof of Proposition 2
Step 1. To prove that the sample paths of the Gaussian field $W_{\Phi}$ are almost surely continuous on $\mathbb{R}\times(0,\infty)$, according to [15, p.73 and p.93], it suffices to show the existence of a constant $q>0$ such that for any $\delta>0$, there exists a constant $C_{\delta}$ satisfying
@@
To prove that sample paths of $W_{\Phi}$ are continuously differentiable almost surely, according to [38], it suffices to show that (a) the difference quotient of $W_{\Phi}$ along the time direction converges to $\partial_{t}W_{\Phi}$ in quadratic mean, (b) the difference quotient of $W_{\Phi}$ along the scale direction converges to $\partial_{s}W_{\Phi}$ in quadratic mean, and (c) both $\partial_{t}W_{\Phi}$ and $\partial_{s}W_{\Phi}$ are continuous in quadratic mean and almost surely have continuous realizations. We verify these conditions as follows.
@@
Proof of (c): To prove that $\partial_{t}W_{\Phi}$ and $\partial_{s}W_{\Phi}$ not only are continuous in the quadratic mean but also almost surely have modification that is sample continuous, according to [15, p.73 and p.93] (see also [37, Corollary 4.6]), it suffices to show that for any $\delta>0$, there exists a constant $C_{\delta}$ such that
@@ 7.3 Proof of Corollary 1
Proof of (11): Based on the continuity of the sample paths of $W_{\Phi}$ on $\mathbb{R}\times(0,\infty)$, it suffices to show that the sample paths of the Gaussian field $W_{\Phi}$ are almost surely continuous at the boundary $\mathbb{R}\times\{0\}$ if we define $W_{\Phi}(t,0)=0$ for any $t\in\mathbb{R}$. According to [15, p.73 and p.93], it suffices to show the existence of a constant $q>0$ such that there exists a constant $C$ satisfying
@@ 7.4 Proof of Theorem 1
where the second equality stems from Proposition 1(c). Therefore, for any $s>0$,
@@
where $\|\mathbf{z}\|^{2}=\mathbf{z}^{\top}\mathbf{z}$. By the result $\textup{Cov}\left(S_{\Phi}(s_{1}),S_{\Phi}(s_{2})\right)\geq 0$ given in Proposition 1(c), the Cauchy-Schwarz inequality, the condition in Assumption 1 that the spectral density $p$ is positive almost everywhere, and (72), we have
@@
    For $s$ satisfying $\textup{Cov}(S_{\Phi}(s),S_{\Phi}(s_{1}))=0$, by Proposition 1(c), we have $\mathbb{E}\left[W^{R}_{\Phi}(s)W^{R}_{\Phi}(s_{1})\right]=0$ and $\mathbb{E}\left[W^{R}_{\Phi}(s)W^{I}_{\Phi}(s_{1})\right]=0,$ which imply that $\Gamma(s,s_{1})=\mathbf{0}_{2\times 2}$ and $Q(\mathbf{z};s)=\mathbf{W}_{\perp}^{\top}(s)\mathbf{W}_{\perp}(s)$ for any $\mathbf{z}\in\mathbb{R}^{2}$. This indicates that $\{(\mathbf{z},Q(\mathbf{z},s))\mid\mathbf{z}\in\mathbb{R}^{2}\}$ is a horizontal plane.
By the fact that the supremum of a family of convex functions is still convex, both $M_{1}$ and $M_{2}$ are convex functions defined on $\mathbb{R}^{2}$. According to Alexandrov’s theorem [41, Theorem 2.1], $M_{1}$ and $M_{2}$ have the second derivatives almost everywhere. More precisely, there exists a set $D\subset\mathbb{R}^{2}$, whose complement has Lebesgue measure zero, such that for $k\in\{1,2\}$ and every point $\mathbf{z}\in D$, there is a quadratic expansion in the form
@@
where $\mathbf{v}_{k}(\mathbf{z})$ is the gradient, and $\mathbf{H}_{k}(\mathbf{z})$ is the Hessian matrix of $M_{k}$ at $\mathbf{z}$. The function $o(\|\mathbf{w}\|^{2})$ satisfies $o(\|\mathbf{w}\|^{2})\|\mathbf{w}\|^{-2}\rightarrow 0$ as $\|\mathbf{w}\|\rightarrow 0$. By (75) (see also [40, Example 13.10]),
@@ 7.5 Proof of Theorem 2
The continuity of $S_{Y}(t,s;\omega)$ and the compactness of $0,\mathcal{S}(\omega)]$ allow us to apply Berge’s maximum theorem [21, Theorem 3.4] (see also [2, p. 570] or Lemma 5 in the appendix), implying that the set-valued function $s_{Y}(\cdot;\omega)$ is upper hemicontinuous at every time point $t_{0}\in\mathcal{M}$. ∎
@@ 7.6 Proof of Corollary 2
Using the notation introduced at the beginning of Section 7.5, Theorems 1 and 2 show that there exists a subset $\widetilde{\Omega}$ of $\Omega$ with $\mathbb{P}(\widetilde{\Omega})=1$ such that for any $\omega\in\widetilde{\Omega}$,
@@ 7.7 Proof of Theorem 3
Because the wavelet $\psi$ is assumed to be analytic, $\mathbf{J}$ is a circularly symmetric complex Gaussian random vector, as shown in Proposition 1(a). By Lemma 1, for any $u>\mathbb{E}[g(\mathbf{J})]$,
@@ 7.9 Proof of Lemma 1
By the Gaussian concentration inequality [23, Corollary 8.5], for any $u>\mathbb{E}[g_{2d}(\mathbf{A}_{2d}\mathbf{N}_{2d})]$,
@@ 7.10 Proof of Lemma 2
Under Assumption 3(b1), $p(\lambda)=L(|\lambda|^{-1})|\lambda|^{H-1}$, where $L$ is slowly varying at infinity. According to [4, Theorem 1.5.3], for any slowly varying function $L$ and any $\delta>0$, $L(s)s^{-\delta}\rightarrow 0$ as $s\rightarrow\infty$. Thus, (115) implies that
@@ 7.11 Proof of Lemma 3
The proof idea can be traced back to the chaining argument in [45]. We firstly create a sequence of sets, denoted as $\{\pi_{j}\}_{j\in\mathbb{N}\cup\{0\}}$, of finite points in $0,\infty)$ such that for each $j\in\mathbb{N}\cup\{0\}$,
@@
By Proposition 1(b), the random variable
@@ 7.12 Proof of Lemma 4
[17, Lemma (4.3.16)] If $\mathbb{X}$ is a topological space, $\mathbb{Y}$ a complete metric space and $r:\mathbb{D}\mapsto\mathbb{Y}$ a continuous mapping defined on a dense subset $\mathbb{D}$ of the space $\mathbb{X}$, then the mapping $r$ is extendable to a continuous mapping defined on the set consisting of all points of $\mathbb{X}$ at which the oscillation of $r$ is equal to zero. Here, we say that the oscillation of the mapping $r$ at a point $x\in\mathbb{X}$ is equal to zero if for every $\varepsilon>0$ there exists a neighborhood $U$ of the point $x$ such that